# How to Evaluate Medical AI

### v1

@@ Abstract
We evaluate our approach using 360 medical dialogues, comparing multiple large language models (LLMs) against a panel of physicians. Large-scale study shows that top-performing models, such as DeepSeek-V3, achieve consistency on par with or exceeding expert consensus. Moreover, we demonstrate that expert judgments exhibit significant variability — often greater than that between AI and the humans. This finding underscores the limitations of any ‘absolute’ metrics and supports the need to adopt relative metrics in medical AI.
@@
Ilia Kopanichuk1, 2, Petr Anokhin1, Vladimir Shaposhnikov1, 3, Vladimir Makharev1, 4, Ekaterina Tsapieva5, Iaroslav Bespalov1, Dmitry V. Dylov1,3, Ivan Oseledets1,3 1AIRI, Moscow, Russia 2Moscow Institute of Physics and Technology, Dolgoprudny, Russia 3Skokovo Institute of Science and Technology, Moscow, Russia 4Innopolis University, Innopolis, Russia 5SberHealth, Moscow, Russia Correspondence: kopanichuk@airi.net
@@ 4.4 Cohen’s Kappa
  • Top-1 only. Standard $\kappa$ considers only the first-ranked diagnosis, ignoring the relevance of other hypotheses, which makes it ill-suited to multi-choice settings.
  • Top-$k$ bias. Generalizing $\kappa$ to top-$k$ metrics triggers the "inflated agreement" effect Carpentier et al. (2017), where response aggregation artificially boosts consensus estimates while obscuring meaningful variation.
@@
In contrast, the RPAD and RRAD distributions computed from the same unbiased estimates but normalized by inter-expert variance—remain well separated and robust to the choice of reference expert and number of hypotheses. Therefore, under high variability in expert opinions, the relative metrics (RPAD, RRAD) provide a more reliable and interpretable means of comparing algorithmic diagnostic performance.

### v2

@@ Abstract
We evaluate our approach using 360 medical dialogues, comparing multiple large language models (LLMs) against a panel of physicians. Large-scale study shows that top-performing models, such as DeepSeek-V3, achieve consistency on par with or exceeding expert consensus. Moreover, we demonstrate that expert judgments exhibit significant variability - often greater than that between AI and the humans. This finding underscores the limitations of any absolute metrics and supports the need to adopt relative metrics in medical AI.
@@
Ilia Kopanichuk1, 2, Petr Anokhin1, 6, Vladimir Shaposhnikov1, 3, Vladimir Makharev1, 4, Ekaterina Tsapieva5, Iaroslav Bespalov1, Dmitry V. Dylov1,3, Ivan Oseledets1,3 1AIRI, Moscow, Russia 2Moscow Institute of Physics and Technology, Dolgoprudny, Russia 3Skokovo Institute of Science and Technology, Moscow, Russia 4Innopolis University, Innopolis, Russia 5SberHealth, Moscow, Russia 6Lomonosov Moscow State University, Moscow, Russia Correspondence: kopanichuk@airi.net
@@ 4.4 Cohen’s Kappa
  • Top‑1 only. Standard $\kappa$ considers only the first‐ranked diagnosis, ignoring the relevance of other hypotheses, which makes it ill‐suited to multi‑choice settings.
  • Top‑$k$ bias. Generalizing $\kappa$ to top‑$k$ metrics triggers the "inflated agreement" effect Carpentier et al. (2017), where response aggregation artificially boosts consensus estimates while obscuring meaningful variation.
@@
In contrast, the RPAD and RRAD distributions computed from the same unbiased estimates but normalized by inter‑expert variance—remain well separated and robust to the choice of reference expert and number of hypotheses. Therefore, under high variability in expert opinions, the relative metrics (RPAD, RRAD) provide a more reliable and interpretable means of comparing algorithmic diagnostic performance.